A set of normally distributed data has a mean of 3.2 and a standard deviation of 0.7. Find the probability of randomly selecting 30 values and getting a mean greater than 3.6.
Answer by stanbon(75887)

A set of normally distributed data has a mean of 3.2 and a standard deviation of 0.7. Find the probability of randomly selecting 30 values and getting a mean greater than 3.6.
-------
Note: Standard deviation for the mean of the sample means = s/sqrt(n)
-------
Your Problem::
z(3.6) = (3.6-3.2)/[0.7/sqrt(30)] = 0.4*sqrt(30)/0.7 = 3.1300
---
P(x-bar > 3.6) = normalcdf(3.13,100) = 0.0008750
===============
Cheers,
Stan H.